# A Covariant and Manifestly Projective Invariant Formulation of Thomas-Whitehead Gravity

### v1

@@ 2.4 Manifestly Projective Invariant form of the TW Field Equations
The equations of motion with respect to the independent fields ${\mathcal{D}}_{ab}$, $\Pi^{c}{}_{ab}$, and $g^{mn}$ were originally derived in [20]. The field equations for $\mathcal{D}_{ab}$ are  [^5] We have divided by a factor of $2\sqrt{|g|}$ from the original version presented in [20]. This makes the equation tensorial, though it is not manifest here.
$\mathbb{D}^{ab}\equiv-\tfrac{({\rm d}-1)}{4\kappa_{0}J_{0}}g^{ab}+\tfrac{1}{2}\hat{\mathcal{K}}_{c}^{\ (ab)c}+\tfrac{\lambda_{0}^{2}}{\sqrt{|g|}}\breve{\nabla}_{c}(\sqrt{|g|}K^{(ab)c})-\lambda_{0}^{2}g_{c}\breve{\mathcal{K}}^{(ab)c}=0$ (2.59)
@@
$\hat{\mathcal{K}}_{d}^{\ abc}=$ $\mathcal{K}^{\bar{d}}_{\ \bar{a}\bar{b}\bar{c}}\mathcal{G}_{d\bar{d}}^{\ \ \ a\bar{a}[b|\bar{b}|c]\bar{c}}$ (2.60)
$\mathcal{G}_{d\bar{d}}^{\ \ \ a\bar{a}b\bar{b}c\bar{c}}=$ $G_{d\bar{d}}g^{a\bar{a}}g^{b\bar{b}}g^{c\bar{c}}-4\delta^{b}_{\ d}\delta^{\bar{b}}_{\ \bar{d}}g^{a\bar{a}}g^{c\bar{c}}+\delta^{b}_{\ d}\delta^{\bar{b}}_{\ \bar{d}}g^{ac}g^{\bar{a}\bar{c}}$ (2.61)
$\breve{\mathcal{K}}^{abc}=$ $g^{ap}g^{bm}g^{cn}\breve{\nabla}_{[m}{\mathcal{D}}_{n]p}~{}~{}~{}.$ (2.62)
The field equations for $\Pi^{c}{}_{ab}$ are  [^6] We have divided by a factor of $2\sqrt{|g|}$ from the original version presented in [20].
~$\tilde{\mathcal{E}}_{c}{}^{ab}\equiv\mathcal{E}_{c}^{\ ab}-\tfrac{1}{d+1}\delta_{c}^{\ (a}\mathcal{E}_{d}^{\ b)d}=0$ (2.63)
$\mathcal{E}_{c}^{\ ab}=\mathcal{E}_{c}^{\ ba}\equiv\tfrac{1}{4\kappa_{0}J_{0}\sqrt{|g|}}\breve{\nabla}_{c}(\sqrt{|g|}g^{ab})-\tfrac{1}{2\sqrt{|g|}}\breve{\nabla}_{d}(\sqrt{|g|}\hat{\mathcal{K}}_{c}^{\ (ab)d})\quad$
$\qquad\qquad\qquad\qquad+\tfrac{\lambda_{0}^{2}}{\sqrt{|g|}}\breve{\nabla}_{d}(\sqrt{|g|}g_{c}\breve{\mathcal{K}}^{(ab)d})-\lambda_{0}^{2}K^{(ab)d}\mathcal{D}_{dc}~{}~{}~{}.$ (2.64)
The field equations for $g^{mn}$ without any external matter sources are  [^7] This equation is equivalent to that originally presented in [20], though written in a slightly different way such that each individual term is projectively invariant.
$\mathcal{H}_{mn}=$ $-\mathcal{K}_{mn}-\kappa_{0}g_{mn}\left[\tfrac{2J_{0}\lambda_{0}^{2}}{({\rm d}+1)\sqrt{|g|}}\breve{\nabla}_{a}(\sqrt{|g|}{\mathcal{K}}^{a}{}_{bcd}K^{bcd})+\mathcal{L}\right]$ (2.65)
$+2\kappa_{0}J_{0}\lambda_{0}^{2}(K_{mcd}K_{n}{}^{cd}+2K_{bcm}K^{bc}{}_{n})+2\kappa_{0}J_{0}(8{\mathcal{K}}_{mb}{\mathcal{K}}^{b}{}_{n}-2{\mathcal{K}}{\mathcal{K}}_{mn})$ (2.66)
$+2\kappa_{0}J_{0}({\mathcal{K}}_{mbcd}{\mathcal{K}}_{n}{}^{bcd}-{\mathcal{K}}^{a}{}_{mcd}{\mathcal{K}}_{an}{}^{cd}-2{\mathcal{K}}^{abc}{}_{m}{\mathcal{K}}_{abcn})$ (2.67)
[…]
$\mathcal{L}=$ $-\tfrac{1}{2\kappa_{0}}{\mathcal{K}}+J_{0}\lambda_{0}^{2}K_{bcd}K^{bcd}-J_{0}(\mathcal{K}_{abcd}\mathcal{K}^{abcd}-4\mathcal{K}_{bd}\mathcal{K}^{bd}+\mathcal{K}^{2})~{}~{}~{}.$ (2.69)
All equations of motion are manifestly projectively invariant as they are composed entirely of projectively invariant objects $\mathcal{K}^{a}{}_{bcd}$, ˘$\breve{\nabla}_{a}$, $g_{ab}$, $g_{a}$, and $\mathcal{D}_{ab}$. None are manifestly tensorial, and in fact $\mathcal{E}_{c}{}^{ab}$ is not tensorial, though $\mathbb{D}^{ab}$ and $\mathcal{H}_{mn}$ are tensorial.
2.5 Covariant Form of the TW Field Equations
@@
$\mathbb{D}^{ab}=$ $-\tfrac{d-1}{4J_{0}\kappa_{0}}g^{ab}-(d-1)g^{ab}\mathcal{K}+2(2d-3)\mathcal{K}^{ab}+\lambda_{0}^{2}(\nabla_{c}-({\rm d}+2)\text{\textcent}_{c})K^{(ab)c}$ (2.70)
which is manifestly tensorial for the reasons described above. To illustrate the tensorial nature of $\mathbb{H}_{mn}$, we express it in the familiar form of the Einstein Equations
$\hat{R}_{mn}-\tfrac{1}{2}g_{mn}\hat{R}=\kappa_{0}\hat{T}_{mn}$ (2.71)
$\hat{T}_{mn}=-\frac{1}{\kappa_{0}}\Big{(}Q_{mn}-\tfrac{1}{2}g_{mn}Q+({\rm d}-1)\mathcal{P}_{mn}-{\rm d}{\mathcal{P}}_{[mn]}-\tfrac{{\rm d}-1}{2}g_{mn}\mathcal{P}\Big{)}$
$\qquad+J_{0}g_{mn}\left[2\lambda_{0}^{2}(-\tfrac{1}{{\rm d}+1}\nabla_{a}+\text{\textcent}_{a})K_{bcd}\mathcal{K}^{abcd}-\lambda_{0}^{2}K^{bcd}K_{bcd}\right.$ (2.72)
$\left.\qquad\qquad\qquad+\mathcal{K}_{abcd}\mathcal{K}^{abcd}-4\mathcal{K}^{ab}\mathcal{K}_{ab}+\mathcal{K}^{2}\right]+2J_{0}\lambda_{0}^{2}(K_{mcd}K_{n}{}^{cd}+2K_{bcm}K^{bc}{}_{n})$
$\qquad+2J_{0}(8{\mathcal{K}}_{mb}{\mathcal{K}}^{b}{}_{n}-2{\mathcal{K}}{\mathcal{K}}_{mn})+2J_{0}({\mathcal{K}}_{mbcd}{\mathcal{K}}_{n}{}^{bcd}-{\mathcal{K}}^{a}{}_{mcd}{\mathcal{K}}_{an}{}^{cd}-2{\mathcal{K}}^{abc}{}_{m}{\mathcal{K}}_{abcn})$
where recall ${\mathcal{P}}_{mn}$, ${\mathcal{P}}$, $Q_{mn}$, and $Q$ were defined in Eqs. (2.17), (2.34), (2.36), and (2.56).
@@
$\mathcal{E}_{c}{}^{ab}=$ $\mathbb{E}_{c}{}^{ab}+g_{c}\mathbb{D}^{ab}+\delta^{(a}{}_{c}\mathfrak{g}^{b)}$ (2.73)
where for brevity we have defined
$\mathbb{E}_{c}{}^{ab}\equiv$ $(\nabla_{c}-2\text{\textcent}_{c})\mathscr{K}^{ab}-\delta^{(a}{}_{c}\text{\textcent}_{d}\mathscr{K}^{b)d}-[(\nabla_{d}-({\rm d}+1)\text{\textcent}_{d})\mathcal{K}_{c}{}^{(ab)d}-2\text{\textcent}_{c}\mathcal{K}^{ab}]$ (2.74)
$-\lambda_{0}^{2}(\mathcal{P}_{cd}-\nabla_{c}\text{\textcent}_{d}-\text{\textcent}_{c}\text{\textcent}_{d})K^{(ab)d}$ (2.75)
$\mathscr{K}^{ab}\equiv$ $(\tfrac{1}{4J_{0}\kappa_{0}}+\mathcal{K})g^{ab}-4\mathcal{K}^{ab}$ (2.76)
$\mathfrak{g}^{b}\equiv$ $\tfrac{1}{4J_{0}\kappa_{0}}g_{d}g^{bd}-\tfrac{1}{2}(\nabla_{d}-({\rm d}+3)\text{\textcent}_{d})(g^{bd}\mathcal{K}-4\mathcal{K}^{bd})~{}~{}~{}.$ (2.77)
The non-tensorality of $\mathcal{E}_{c}{}^{ab}$ is explicitly seen in the second and third terms of Eq. (2.73) along with the first term of Eq. (2.77): these take the form of the non-tensorial $g_{c}$ multiplying a tensor, the product of which is necessarily non-tensorial. An obvious question is: does the actual equation of motion $\tilde{E}_{c}{}^{ab}$, formed by subtracting from $\mathcal{E}_{c}{}^{ab}$ its trace, turn out to be tensorial? The answer is no as seen below
~$\tilde{\mathcal{E}}_{c}{}^{ab}=$ $\mathbb{E}_{c}{}^{ab}-\tfrac{1}{{\rm d}+1}\delta^{(a}{}_{c}\mathbb{E}_{m}{}^{b)m}+g_{c}\mathbb{D}^{ab}-\tfrac{1}{{\rm d}+1}\delta^{(a}{}_{c}g_{m}\mathbb{D}^{b)m}$ (2.78)
@@
Regardless, making the tensorial or non-tensorial nature of the equations manifest has now completely hidden their projective invariance. Now written in terms of the manifestly tensorial ${\mathcal{P}}_{ab}$, $\nabla_{a}$, ¢$\text{\textcent}_{a}$, and $Q_{ab}$, we have to expose the projective transformations on each of these, and show that they cancel, to demonstrate that the equations are indeed still projectively invariant. Under a projective transformation, Eq. (2.1), these transform as
$\mathcal{P}_{ab}\to$ $\mathcal{P}_{ab}+\nabla_{a}v_{b}-v_{a}v_{b}$ (2.79)
¢$\text{\textcent}_{a}\to$$\text{\textcent}_{a}+v_{b}$ (2.80)
$Q_{mn}\to$ $Q_{mn}+\nabla_{m}v_{n}-{\rm d}\nabla_{n}v_{m}+({\rm d}-1)v_{m}v_{n}~{}~{}~{}.$ (2.81)
@@
$\mathcal{P}_{cd}-\nabla_{c}\text{\textcent}_{d}-\text{\textcent}_{c}\text{\textcent}_{d}~{}~{}~{},~{}~{}~{}Q_{mn}-\tfrac{1}{2}g_{mn}Q+({\rm d}-1)\mathcal{P}_{mn}-{\rm d}{\mathcal{P}}_{[mn]}-\tfrac{{\rm d}-1}{2}g_{mn}\mathcal{P}$ (2.82)
$(-\tfrac{1}{{\rm d}+1}\nabla_{a}+\text{\textcent}_{a})K_{bcd}\mathcal{K}^{abcd}~{}~{}~{},~{}~{}~{}(\nabla_{c}-2\text{\textcent}_{c})\mathscr{K}^{ab}-\delta^{(a}{}_{c}\text{\textcent}_{d}\mathscr{K}^{b)d}$ (2.83)
$[(\nabla_{d}-({\rm d}+1)\text{\textcent}_{d})\mathcal{K}_{c}{}^{(ab)d}-2\text{\textcent}_{c}\mathcal{K}^{ab}]~{}~{}~{},~{}~{}~{}(\nabla_{c}-({\rm d}+2)\text{\textcent}_{c})K^{(ab)c}$ (2.84)
$(\nabla_{d}-({\rm d}+3)\text{\textcent}_{d})(g^{bd}\mathcal{K}-4\mathcal{K}^{bd})~{}~{}~{}.$ (2.85)
This section and the previous section have served to illustrate the main driving force for the introduction of manifestly tensorial and projectively invariant fields to which we now turn.
3 Covariant and Manifestly Projective Invariant
Formulation of TW Gravity
3.1 Projectively Invariant Connection and Schouten Tensor

### v3

@@ II.4 Manifestly Projective Invariant form of the TW Field Equations
We note that from here on, we work in units where the speed of light $c=1$, but keep the constants $J_{0}$, $\kappa_{0}$, and $\lambda_{0}$ owing to the discussion in the previous section. The equations of motion with respect to the independent fields ${\mathcal{D}}_{ab}$, $\Pi^{c}{}_{ab}$, and $g^{mn}$ were originally derived in [18] in constant-volume coordinates. Here we review the more general, non-constant volume coordinate version of the field equations originally derived in [20, 21]. The field equations for $\mathcal{D}_{ab}$ are  [^4] We have divided by a factor of $2\sqrt{|g|}$ from the original version presented in [20, 21]. This makes the equation tensorial, though it is not manifest here.
$0=\mathbb{D}^{ab}\equiv$ $-\tfrac{({\rm d}-1)}{4J_{0}\kappa_{0}}g^{ab}+\tfrac{1}{2}\hat{\mathcal{K}}_{c}^{\ (ab)c}$ (69)
$+\tfrac{\lambda_{0}^{2}}{\sqrt{|g|}}\breve{\nabla}_{c}(\sqrt{|g|}K^{(ab)c})-\lambda_{0}^{2}g_{c}\breve{\mathcal{K}}^{(ab)c}$ (70)
@@
$\hat{\mathcal{K}}_{d}^{\ abc}=\mathcal{K}^{\bar{d}}_{\ \bar{a}\bar{b}\bar{c}}\mathcal{G}_{d\bar{d}}^{\ \ a\bar{a}[b|\bar{b}|c]\bar{c}}$ (71)
$\mathcal{G}_{d\bar{d}}^{\ \ \ a\bar{a}b\bar{b}c\bar{c}}$
$\quad=G_{d\bar{d}}g^{a\bar{a}}g^{b\bar{b}}g^{c\bar{c}}-4\delta^{b}_{\ d}\delta^{\bar{b}}_{\ \bar{d}}g^{a\bar{a}}g^{c\bar{c}}+\delta^{b}_{\ d}\delta^{\bar{b}}_{\ \bar{d}}g^{ac}g^{\bar{a}\bar{c}}$ (72)
$\breve{\mathcal{K}}^{abc}=g^{ap}g^{bm}g^{cn}\breve{\nabla}_{[m}{\mathcal{D}}_{n]p}~{}.$ (73)
Note that $\hat{{\mathcal{K}}}_{d}^{\ abc}$, that includes $G_{d\bar{d}}$ in its definition, should not be confused with ${\mathcal{K}}_{d}^{\ abc}=g_{dm}{\mathcal{K}}^{mabc}$ that, here on $\mathcal{M}$, has its indices raised or lowered via $g^{dm}$ or $g_{dm}$ as in Eq. (58).
The field equations for $\Pi^{c}{}_{ab}$ are  [^5] We have divided by a factor of $2\sqrt{|g|}$ from the original version presented in [20, 21].
$0=\tilde{\mathcal{E}}_{c}{}^{ab}\equiv$ $\mathcal{E}_{c}^{\ ab}-\tfrac{1}{d+1}\delta_{c}^{\ (a}\mathcal{E}_{d}^{\ b)d}$ (74)
$\mathcal{E}_{c}^{\ ab}\equiv$ $\tfrac{1}{4J_{0}\kappa_{0}\sqrt{|g|}}\breve{\nabla}_{c}(\sqrt{|g|}g^{ab})-\tfrac{1}{2\sqrt{|g|}}\breve{\nabla}_{d}(\sqrt{|g|}\hat{\mathcal{K}}_{c}^{\ (ab)d})$
$+\tfrac{\lambda_{0}^{2}}{\sqrt{|g|}}\breve{\nabla}_{d}(\sqrt{|g|}g_{c}\breve{\mathcal{K}}^{(ab)d})-\lambda_{0}^{2}K^{(ab)d}\mathcal{D}_{dc}.$ (75)
The field equations for $g^{mn}$ without any external matter sources are  [^6] This equation is equivalent, up to the addition of the cosmological constant $\Lambda_{0}$, to that originally presented in [20, 21], though written in a slightly different way such that each individual term is projectively invariant.
$0=\mathcal{H}_{mn}\equiv-\mathcal{K}_{mn}+\kappa_{0}\left(2\mathbb{L}_{mn}-g_{mn}\mathcal{L}\right)$ (76)
$\qquad\qquad\quad-g_{mn}\tfrac{2J_{0}\kappa_{0}\lambda_{0}^{2}}{({\rm d}+1)\sqrt{|g|}}\breve{\nabla}_{a}(\sqrt{|g|}{\mathcal{K}}^{a}{}_{bcd}K^{bcd})$ (77)
$\mathbb{L}_{mn}=J_{0}\lambda_{0}^{2}(K_{mcd}K_{n}^{\ cd}+2K_{bcm}K^{bc}_{\ \ n})$ (78)
$\qquad\quad+J_{0}({\mathcal{K}}_{mbcd}{\mathcal{K}}_{n}^{\ bcd}-{\mathcal{K}}^{a}_{\ mcd}{\mathcal{K}}_{an}^{\ \ cd}-2{\mathcal{K}}^{abc}_{\ \ \ m}{\mathcal{K}}_{abcn})$ (79)
$\qquad\quad+J_{0}\left(8{\mathcal{K}}_{mb}{\mathcal{K}}^{b}_{\ n}-2{\mathcal{K}}{\mathcal{K}}_{mn}\right)$ (80)
where the TW Lagrangian $\mathcal{L}$ and PGB Lagrangian $\mathbb{L}$ were defined in Eqs. (54) and (55), respectively. Recall all indices are raised or lowered via $g^{dm}$ or $g_{dm}$ here on $\mathcal{M}$ as explained in Eqs. (57), (58), and (59).
All equations of motion are manifestly projective invariant as they are composed entirely of projectively invariant objects $\mathcal{K}^{a}{}_{bcd}$, ˘$\breve{\nabla}_{a}$, $g_{ab}$, $g_{a}$, and $\mathcal{D}_{ab}$. None are covariant, and in fact $\mathcal{E}_{c}{}^{ab}$ is not tensorial, though $\mathbb{D}^{ab}$ and $\mathcal{H}_{mn}$ are tensorial as illustrated in the next section.
II.5 Covariant Form of the TW Field Equations
@@
$0=\mathbb{D}^{ab}=\lambda_{0}^{2}(\nabla_{c}-({\rm d}+2)\text{\textcent}_{c})K^{(ab)c}-({\rm d}-1)\mathbb{K}^{ab}-2\mathcal{K}^{ab}$ (81)
$\hskip 6.0pt\mathbb{K}^{ab}\equiv g^{ab}\left({\mathcal{K}}+\tfrac{1}{4J_{0}\kappa_{0}}\right)-4\mathcal{K}^{ab}$ (82)
which is covariant for the reasons described above. To illustrate the tensorial nature of $\mathcal{H}_{mn}$, we express it in the familiar form of the Einstein Equations
$\hat{R}_{mn}-$ $\tfrac{1}{2}g_{mn}\hat{R}=\kappa_{0}\hat{T}_{mn}+g_{mn}\Lambda_{0}$ (83)
$\hat{T}_{mn}=$ $-\kappa_{0}^{-1}\left(Q_{mn}-\tfrac{1}{2}g_{mn}Q\right)+2\mathbb{L}_{mn}-g_{mn}\mathbb{L}$ (84)
$-\kappa_{0}^{-1}\left(({\rm d}-1)\mathcal{P}_{mn}-{\rm d}{\mathcal{P}}_{[mn]}-\tfrac{{\rm d}-1}{2}g_{mn}\mathcal{P}\right)$ (85)
$-g_{mn}\tfrac{2J_{0}\lambda_{0}^{2}}{{\rm d}+1}(\nabla_{a}-({\rm d}+1)\text{\textcent}_{a})(K_{bcd}\mathcal{K}^{abcd})$ (86)
where recall ${\mathcal{P}}_{mn}$, ${\mathcal{P}}$, $Q_{mn}$, $Q$, $\mathbb{L}_{mn}$, and $\mathbb{L}$ were defined in Eqs. (20), (40), (42), (65), (66), (78), and (55).
@@
$\mathcal{E}_{c}{}^{ab}=$ $\mathbb{E}_{c}{}^{ab}+g_{c}\mathbb{D}^{ab}+\delta^{(a}{}_{c}\mathfrak{g}^{b)}$ (87)
where for brevity we have defined the tensor $\mathbb{E}_{c}^{\ ab}$ and the non-tensorial object $\mathfrak{g}^{b}$ as
$\mathbb{E}_{c}{}^{ab}\equiv$ $(\nabla_{c}-2\text{\textcent}_{c})\mathbb{K}^{ab}-\delta^{(a}{}_{c}\text{\textcent}_{d}\mathbb{K}^{b)d}$ (88)
$-[(\nabla_{d}-({\rm d}+1)\text{\textcent}_{d})\mathcal{K}_{c}{}^{(ab)d}-2\text{\textcent}_{c}\mathcal{K}^{ab}]$ (89)
$-\lambda_{0}^{2}(\mathcal{P}_{cd}-\nabla_{c}\text{\textcent}_{d}-\text{\textcent}_{c}\text{\textcent}_{d})K^{(ab)d}$ (90)
$\mathfrak{g}^{b}\equiv$ $\tfrac{1}{4J_{0}\kappa_{0}}g_{d}g^{bd}-\tfrac{1}{2}(\nabla_{d}-({\rm d}+3)\text{\textcent}_{d})(g^{bd}\mathcal{K}-4\mathcal{K}^{bd}).~{}~{}~{}~{}~{}~{}$ (91)
Recall here on $\mathcal{M}$ the tensor ${\mathcal{K}}_{c}^{\ abd}$ has its indices raised or lowered via $g^{cm}$ or $g_{cm}$ as in Eqs. (57) and (58).
The non-tensorality of $\mathcal{E}_{c}{}^{ab}$ is explicitly seen in the second and third terms of Eq. (87) along with the first term of Eq. (91): these take the form of the non-tensorial $g_{c}$ multiplying a tensor, the product of which is necessarily non-tensorial. An obvious question is: does the actual equation of motion $\tilde{E}_{c}{}^{ab}$, formed by subtracting from $\mathcal{E}_{c}{}^{ab}$ its trace, turn out to be tensorial? The answer is no as seen below
$0=\tilde{\mathcal{E}}_{c}^{\ ab}=$ $\tilde{\mathbb{E}}_{c}{}^{ab}+g_{c}\mathbb{D}^{ab}-\tfrac{1}{{\rm d}+1}\delta^{(a}_{\ \ c}g_{m}\mathbb{D}^{b)m}$ (92)
$\tilde{\mathbb{E}}_{c}{}^{ab}\equiv$ $\mathbb{E}_{c}{}^{ab}-\tfrac{1}{{\rm d}+1}\delta_{c}{}^{(a}\mathbb{E}_{m}{}^{b)m}$ (93)
@@
Regardless, making the tensorial or non-tensorial nature of the equations manifest has now completely hidden their projective invariance. Now written in terms of the covariant ${\mathcal{P}}_{ab}$, $\nabla_{a}$, ¢$\text{\textcent}_{a}$, and $Q_{ab}$, we have to expose the projective transformations on each of these, and show that they cancel, to demonstrate that the equations are indeed still projectively invariant. Under a projective transformation, Eq. (1), these transform as
$\mathcal{P}_{ab}\to$ $\mathcal{P}_{ab}+\nabla_{a}v_{b}-v_{a}v_{b}$ (94)
¢$\text{\textcent}_{a}\to$$\text{\textcent}_{a}+v_{b}$ (95)
$Q_{mn}\to$ $Q_{mn}+\nabla_{m}v_{n}-{\rm d}\nabla_{n}v_{m}+({\rm d}-1)v_{m}v_{n}~{}.$ (96)
@@
$\mathcal{P}_{cd}-\nabla_{c}\text{\textcent}_{d}-\text{\textcent}_{c}\text{\textcent}_{d}~{}~{}~{}$ (97)
$Q_{mn}-\tfrac{1}{2}g_{mn}Q+({\rm d}-1)\mathcal{P}_{mn}-{\rm d}{\mathcal{P}}_{[mn]}-\tfrac{{\rm d}-1}{2}g_{mn}\mathcal{P}$ (98)
$(-\tfrac{1}{{\rm d}+1}\nabla_{a}+\text{\textcent}_{a})K_{bcd}\mathcal{K}^{abcd}~{}~{}~{}$ (99)
$(\nabla_{c}-2\text{\textcent}_{c})\mathbb{K}^{ab}-\delta^{(a}{}_{c}\text{\textcent}_{d}\mathbb{K}^{b)d}$ (100)
$(\nabla_{d}-({\rm d}+1)\text{\textcent}_{d})\mathcal{K}_{c}{}^{(ab)d}-2\text{\textcent}_{c}\mathcal{K}^{ab}~{}~{}~{}$ (101)
$(\nabla_{c}-({\rm d}+2)\text{\textcent}_{c})K^{(ab)c}$ (102)
$(\nabla_{d}-({\rm d}+3)\text{\textcent}_{d})(g^{bd}\mathcal{K}-4\mathcal{K}^{bd})~{}.$ (103)
This section and the previous section have served to illustrate the main driving force for the introduction of covariant and projectively invariant fields to which we now turn.
III Covariant and Manifestly Projective Invariant Formulation of TW Gravity
III.1 Projectively Invariant Connection and Schouten Tensor